I need help trying to find the difference quotient with this function: 3-5x^2
Answer by stanbon(75887)

I need help trying to find the difference quotient with this function: 3-5x^2
-----------------------------------
f(x) = -3x^2 + 3
f(x+h)=-3(x+h)^2 + 3
f(x+h)=-3x^2-6hx-3h^2+3
------------------------------
f(x+h)-f(x) = -6hx-3h^2
------------
[f(x+h)-f(x)]h = -6x-3h
================================
Cheers,
Stan H.